If zbase1=1−2i,zbase2=−7+i. Findzbase1−zbase2
If z1=1−2i,z2=−7+i. Findz1−z2z1−z2
a.\\(8-3i\\)
b.\\(-7-2i\\)
c.\\(7+2i\\)
d.\\(8-2i\\)
1
Expert's answer
2020-06-04T19:35:13-0400

Given

"z_1 = 1 - 2i" and "z_2 = -7 + i"


"z_1 - z_2 = (1 - 2i) - ( -7 + i ) = 1 - 2i + 7 - i"

"= 8 - 3i"


For next part,

"z_1 - z_2z_1 - z_2 = (1 - 2i) - (-7+i)(1-2i) - (-7+i)"

"= 1 - 2i - (-7 + i + 14i - 2i^{2})+ 7 - i"

"= 8 - 3i +5 - 15i"

"13 - 18i"

where "i^{2} = -1"